Calculate the volume of a 2.5kg block, if its density is 8 kg/m3.

a. 0.31 m3 b. 1.50 m3 c. 3.20 m3 d. 20 m3




1
Expert's answer
2021-04-05T11:13:04-0400
"V=\dfrac{m}{\rho}=\dfrac{2.5\ kg}{8\ \dfrac{kg}{m^3}}=0.31\ m^3."

Answer:

a. 0.31 m3